Reduce to canonical form :

𝑒π‘₯π‘₯ βˆ’ 𝑒π‘₯𝑦 + 𝑒𝑦𝑦 + 𝑒π‘₯ = 0



1
Expert's answer
2022-01-10T15:59:24-0500
"\Delta = a_{12}^2 - {a_{11}}{a_{22}} = {(-\frac{1}{2})^2} - 1 \cdot 1 = - \frac{3}{4} < 0"

Β - we have an equation of elliptic type.

Characteristic equation:


"\frac{{dy}}{{dx}} = \frac{{{a_{12}} + \sqrt \Delta }}{{{a_{11}}}} = \frac{{-\frac{1}{2} + \frac{{\sqrt 3 }}{2}i}}{1} \Rightarrow y =- \frac{x}{2} + \frac{{\sqrt 3 }}{2}ix + {C_1}"


or


"\frac{{dy}}{{dx}} = \frac{{{a_{12}} - \sqrt \Delta }}{{{a_{11}}}} = \frac{{-\frac{1}{2} - \frac{{\sqrt 3 }}{2}i}}{1} \Rightarrow y = -\frac{x}{2} - \frac{{\sqrt 3 }}{2}ix + {C_2}"


Let


"\xi (x,y) = \frac{{{C_1} + {C_2}}}{2}""= \frac{1}{2}\left( {y + \frac{x}{2} - \frac{{\sqrt 3 }}{2}ix + y+\frac{x}{2} + \frac{{\sqrt 3 }}{2}ix} \right) = y +\frac{x}{2}"


"\eta \left( {x,y} \right) = \frac{{{C_2} - {C_1}}}{{2i}}""= \frac{1}{{2i}}\left( {y +\frac{x}{2} + \frac{{\sqrt 3 }}{2}ix - y - \frac{x}{2} + \frac{{\sqrt 3 }}{2}ix} \right) = \frac{{\sqrt 3 }}{2}x"


"u(x,y) = v\left( {\xi ;\,\eta } \right)"


Then


"{u_x} = {v_\xi }{\xi _x} + {v_\eta }{\eta _x} = \frac{1}{2}{v_\xi } + \frac{{\sqrt 3 }}{2}{v_\eta }"




"{u_y} = {v_\xi }{\xi _y} + {v_\eta }{\eta _y} = {v_\xi }"




"{u_{xx}} = \frac{1}{2}\left( {{v_{\xi \xi }}{\xi _x} + {v_{\xi \eta }}{\eta _x}} \right) + \frac{{\sqrt 3 }}{2}\left( {{v_{\eta \xi }}{\xi _x} + {v_{\eta \eta }}{\eta _x}} \right)"




"= \frac{1}{2}\left( { \frac{1}{2}{v_{\xi \xi }} + \frac{{\sqrt 3 }}{2}{v_{\xi \eta }}} \right) + \frac{{\sqrt 3 }}{2}\left( { \frac{1}{2}{v_{\eta \xi }} + \frac{{\sqrt 3 }}{2}{v_{\eta \eta }}} \right)"

"= \frac{1}{4}{v_{\xi \xi }} +\frac{{\sqrt 3 }}{2}{v_{\xi \eta }} + \frac{3}{4}{v_{\eta \eta }}"


"{u_{yy}} = {v_{\xi \xi }}{\xi _y} + {v_{\xi \eta }}{\eta _y} = {v_{\xi \xi }}"




"{u_{yx}} = {v_{\xi \xi }}{\xi _x} + {v_{\xi \eta }}{\eta _x} = \frac{1}{2}{v_{\xi \xi }} + \frac{{\sqrt 3 }}{2}{v_{\xi \eta }}"


Substitute the found values ​​into the original equation:


"\frac{1}{4}{v_{\xi \xi }} +\frac{{\sqrt 3 }}{2}{v_{\xi \eta }} + \frac{3}{4}{v_{\eta \eta }}- \frac{1}{2}{v_{\xi \xi }} -\frac{{\sqrt 3 }}{2}{v_{\xi \eta }}"

"+{v_{\xi \xi }}+\frac{1}{2}{v_\xi } + \frac{{\sqrt 3 }}{2}{v_\eta }=0"


"\frac{3}{4}{v_{\xi \xi }} + \frac{3}{4}{v_{\eta \eta }}=-\frac{1}{2}{v_\xi } - \frac{{\sqrt 3 }}{2}{v_\eta }"


"{v_{\xi \xi }} + {v_{\eta \eta }} =- \frac{2}{3}{v_\xi } - \frac{{2\sqrt 3 }}{3}{v_\eta }"


Answer:Β 

"{v_{\xi \xi }} + {v_{\eta \eta }} = -\frac{2}{3}{v_\xi } - \frac{{2\sqrt 3 }}{3}{v_\eta },"

whereΒ 

"\xi (x,y) = \ y + \frac{x}{2},\ \eta \left( {x,y} \right) = \frac{{\sqrt 3 }}{2}x,"




"u(x,y) = v\left( {\xi ;\,\eta } \right)"